Will Ca(OH)2 precipitate from solution if the pH of a 3.0×10−2 M solution of CaCl2 is adjusted to 8.0?   Will Ag2SO4 precipitate when 100 mL of 4.0×10−2 M AgNO3 is mixed with 10 mL of 5.0×10−2MNa2SO4 solution?
Answer #1

CaCl2 is a strong electrolyte so it will completely dissociate :
CaCl2 -------> Ca(2+) + 2Cl(-)

so concentration of CaCl2 = concentration of Ca(2+) = 0.03 M

pH = 8
so as pH + pOH = 14
pOH = 14-pH = 14-8 = 6
pOH = 6
-log[OH(-)] = 6
log [OH(-)] = -6
[OH(-)] = 10^-6

Ca(OH)2 (s) <--------> Ca(2+) (aq) + 2OH(-) (aq)
concentration of Ca(2+),[Ca(2+)] = 0.03 M
concentration of OH(-),[OH(-)] = 10^-6 M

so Qsp = [Ca(2+)] [OH(-)]^2
Qsp = 0.03 X (10^-6)^2
Qsp = 0.03 X 10^-12 = 3 X 10^-14

from literature Ksp Ca(OH)2 = 6.5 x 10^-6

so as Ksp > Qsp ..Ca(OH)2 will not precipitate

next one also do in the same way

Ans is No silver is not very solubable